Twenty nine less than -10 times a number is equal to -18 times the number plus 91
Answer by Boreal(15235)

number =x
-10x -29= -18x+ 91
8x= 120
x= 15
-10 times that is -150 and 29 less is - 179
-18 times that is - 270 and 91 more than that is -179.
The number is 15.